Bent Gregorian Technical Observing Plan: Jun 2009

Execute TCS startup checklist every night
  • Complete by sunset - 1hr
  • Include mount encoder resets each night.
  • Reset collimation model = clear all active/global offsets
  • At end of startup, issue a standard preset

On-sky TCS SW Tests (B28)
  • B28 - Test 1 - Evaluate new workstations with respect to science software
  • B28 - Test 2 - Evaluate new workstations as a backup for the TO Station
  • B24 - Test 1 - Spherical Motion Test
  • B23-Test 12 - Verify new oacontrol package features with the following changes.
  • B23-Test 13 - Verify that the Tertiary Mirror in the OSS recognizes that a focal station is selected
  • B23-Test 16 - Verify that the all-on-source signal is working
  • B22-Test 1 - Evaluate running GCS as only process on a TCS server
  • B21-Test 3 - Evaluate the updatePointingReference functionality for coordinate types DETXY (pointing origin) and RADEC (on-sky)
  • B17-Test 1 - Evaluate the OffsetPointing functionality with move types both relative and absolute and modes DETXY (pointing origin) and RADEC (on-sky)

Startup pointing & collimation correction (10 minutes):
  • Use WT10 or ACT pointing stars with IRTC
  • Select target at mid EL (50-70 degrees)
  • Preset in PARALLACTIC/TRACK mode:
    With AGw: pointto starname gs=0 Tel="ACQUIRE" PA=0 PAM-
    With IRTC: pointto starname gs=0 Tel="TRACK" PA=0 PAM-
  • If using IRTC, turn on "movie" mode
  • Have TO run "ptspiral" from lbtmu01 and
    • ^C when the star appears on the chip
    • Adjust IE and CA to center; manually focus if necessary
  • Preset in POSITION/ACTIVE mode on axis (Persson std)
  • Allow system to collimate (WFS<400nm rms)

Verify and Update Collimation Model [2hr + reduction + 1 iteration]
  • Prerequisites:
    TELESCOPE IN THERMAL EQUILIBRIUM!
    DECENT SEEING (less than 1.25 arcsec FWHM on guider)
  • Use collimation sources, Persson standards, off-axis guide stars.
  • Set PSF to send collimation updates to M1.
  • Verify M2 is at its nominal position
  • Observe sources: 25(15) < EL < 87 if possible
    1. pointto sourceID gs=0 Tel="ACTIVE"
    2. Allow system to collimate: WFS<400nm rms
    3. takepic 2.0 to record mirror information
    4. Repeat steps 1-3 as needed to cover EL-space
  • Observe minimum 1 full cycle through EL range, 12 sources,
    Prefer two cycles, 20+ sources.
  • Generate new collimation model (J.Hill?)
  • Implement new collimation model

Verify and Update Pointing Model [2hr + reduction + 1 iteration]
  • Prerequisites:
    TELESCOPE IN THERMAL EQUILIBRIUM!
    Using new collimation model from previous step
    Re-measure mechanical rotator center coords on IRTC
    • Take pair of TRACK mode images rotated 180 degrees apart
    • New rotator center: X Y
  • Update collimation (not split, M1 only) at EL 50-70 degrees
  • Update pointing, repeat with ptspiral as needed below
  • Use WT10 and ACT stars
  • Turn on pointing log in PCS GUI
    Filename = ____________________
  • Use /home/LBTO/POINTING/Pointingstars5 to schedule
    1. In your working directory on an observer workstation:
      cp ~LBTO/POINTING/* . to copy program files
    2. Type ./Pointingstars5 at linux prompt to run program
    3. Enter the sequential reference field number [1-88]
    4. Preset to the source listed, TRACK mode, take 10s exposure.
    5. Repeat steps 2-4 for at least 24 points, 48+ would be better.
    6. If Pointingstars5 returns a 999, skip this star and try again later
  • Stop pointing log in PCS GUI
  • Generate new pointing model (D.Thompson)
  • Implement:
    • Copy to ~telescope/TCS/Configuration/PCS/
    • Update LUCIFER entry for LEFTPOINTINGMODEL
      in PCSInstrument.conf
    • Re-authorize LUCIFER on IIF GUI to read in new pointing model
  • Test new pointing model

Optimize Guiding Gain Factor [4 hrs]
  • Minimize guiding residuals
  • Try different values of Gain Factor in the range 0.05-1.0 at, say, ~0.15 step

Demonstrate Guiding/WFSing to R=18 mag [6 hrs]
  • Plot rms vs. magnitude [M92astrom has a good selection of magnitudes]
  • Stone-O Field at R.A.~22hrs ready for late night test (if needed).
  • May require WFS operating to 120s exposures

Lucifer Remote Software test [1hr]
  • Perform a basic test of the communication of the LUCIFER telescope service SW with the TCS (Tue 09th - 1hr before sunrise. Wed 10th backup).

Quantify/Track Technical Downtime and Observing Efficiency

Staffing for Observations

All dates and times in UT, Jun 2009

UTDateUT Instr Observers Start/End/Down Times Comments
07 IRTC DT, MP S: E: D: SW on call for TCS
08 IRTC DT, MP S: E: D:  
09 IRTC DT, MP S: E: D: Walter S. for LUCI at LBT, Volker remotely
10 IRTC DT,MP S: E: D:  
11 IRTC DT, MP S: E: D:  
12 IRTC DT, MP S: E: D:  
13 IRTC DT ,MP S: E: D:  
14 IRTC DT, MP S: E: D:  

MP: Marco Pedani
DT: Dave Thompson
SW: TCS group

-- MarcoPedani - 03 Jun 2009
Topic revision: r5 - 30 Sep 2009, DougMiller
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback